If FEED is coded as 88, then BREAD and BUTTER is coded as ?

Grade:12

1 Answers

Vikas TU
6 years ago
Hi Rahul,
Take this example as: if RED is coded as 6720
As indicated by the request of English Alphabet , 
R = 18 , E = 5 , D = 4 . 
Add 2 to the position of every letters in order and compose the whole in turn around request. 
R = 18+2 =20 , E = 5+2 =7 , D = 4+2 =6 
(20)(7)(6) — > Write backward request. — > (6)(7)(20) = 6720 
Essentially for GREEN. 
GREEN =(7+2)(18+2)(5+2)(5+2)(14+2) = (9)(20)(7)(7)(16). 
Presently switch the request , (16)(7)(7)(20)(9). 
Henceforth the appropriate response is 1677209 .